John can finish a painting job in 2 days. Peter can finish the
job in 3 days. Danny can do the job in 6 days. How long will
it take all of them together to finish the job?
1
Expert's answer
2021-09-17T03:49:02-0400

It follows that John can do "\frac{1}2" part of the job per one day, Peter can do "\frac{1}3" part of the job per one day, and Danny can do "\frac{1}6" part of the job per one day. Taking into accounrt that "\frac{1}{2}+\frac{1}3+\frac{1}6=\frac{3+2+1}{6}=\frac{6}{6}=1,"

we conclude that it will take one day all of them together to finish the job.